given: f(x)=x^2-6x+5
1) write f(x) in standard form
2) find the vertex and x-intercepts of f(x)
3) solve f(x)

Answer by Nate(3500)

f(x)=x^2-6x+5
y = x^2 - 6x + 5
y - 5 = x^2 - 6x
y - 5 + 9 = (x-3)^2
y + 4 = (x-3)^2
f(x) = (x-3)^2 - 4
In standard form y=a(x-h)^+k .... the vertex is (h,k), so it is (3,-4)
For the x-intercepts, y has to equal zero .... so the two answers for x is 5 and 1.
The answers are in a parabola line (U-shaped) starting from (3,-4) and moving upwards on the cartesian coordinate plane.